Let y^2=4ax be a parabola and P be a point on it. Let the normal at P intersect the X-axis at Q. Draw a line at Q perpendicular to the above normal. Show that this line intersects the parabola y^2+4a(x-2a)=0
1
Expert's answer
2020-06-08T20:22:36-0400

Let the co-ordinates of the point P be "(4a,4a)"

Let Slope of tangent at point p is given by "m_1".

"y = \sqrt{4ax}"

"m_1 = \frac{dy}{dx} = \frac{\sqrt{4a}}{2{\sqrt{x}}} = \frac{1}{2}"


Let Slope of normal at P, is "m_2".

"m_2 = -2"


Equation of normal will be

"(y-4a)=-2(x-4a)"

"y = -2x + 12a"


Since this equation will intersect x-axis at Q then co-ordinates of Point Q is "(6a,0)".


Equation of line passing through point Q and perpendicular to PQ will have slope "\frac{1}{2}."

"y- 0 = \frac{1}{2}(x - 6a)"

"y = \frac{1}{2}x - 3a" . . . . . . . . . . .. . . . . . .. .(i)


Equation (i) will intersect parabola "y^{2} + 4a(x-2a) = 0" at "(-2a,-4a)."

This is all is shown in figure.